# 11email: {mai.kassem, mohammad.yaqub}

### v1

@@ Appendix 0.A Appendix
Atelectasis 10
CO (Arterial), HDL, Cholesterol, ELWI (PiCCO), T Low (APRV), GEDI (PiCCO), LDL measured, T High (APRV), LDL calculated, Serum Osmolality
Cardiomegaly 10
BiPap bpm (S/T -Back up), LDL measured, ELWI (PiCCO), D-Dimer, Impaired Skin Length #5, Impaired Skin Width #5, Uric Acid, GEDI (PiCCO), QTc, Cholesterol
Consolidation 9
Manual Blood Pressure Diastolic Right, Manual Blood Pressure Systolic Right, ELWI (PiCCO), GEDI (PiCCO), CFI (PiCCO), Manual Blood Pressure Diastolic Left, Negative Insp. Force, Cholesterol, PCA basal rate (mL/hour)
Edema 10
SV (Arterial), CO (Arterial), ELWI (PiCCO), CFI (PiCCO), GEDI (PiCCO), Bladder Scan Estimate, SVV (Arterial), Gentamicin (Random), LDL measured, BiPap bpm (S/T -Back up)
Enlarged Cardiomediastinum 5
ELWI (PiCCO), GEDI (PiCCO), SVV (Arterial), D-Dimer, RCexp (Measured Time Constant)
Fracture 10
Absolute Count - Monos, CK-MB, Absolute Count - Neuts, Troponin-T, CO2 production, Differential-Bands, Vti High, Absolute Count - Lymphs, Chloride (whole blood), Total Bilirubin
Lung Lesion 10
Temporary Ventricular Sens Setting mV, Temporary Venticular Sens Threshold mV, PCV Level, Absolute Count - Neuts, GI #1 Tube Mark (CM), Temporary Pacemaker Rate, Glucose (whole blood), Ionized Calcium, Total Bilirubin, Absolute Count - Eos
Lung Opacity 6
Cardiac Output (thermodilution), Bladder Scan Estimate, Ammonia, Serum Osmolality, PBP (Prefilter) Replacement Rate, Current Goal
Pleural Effusion 10
SV (Arterial), CO (Arterial), ELWI (PiCCO), Permanent Pacemaker Rate, GEDI (PiCCO), Gentamicin (Random), SVV (Arterial), Arctic Sun/Alsius Temp #2 C, Feeding Weight, Arctic Sun/Alsius Temp #1 C
Pleural Other 10
PCV Level, Impaired Skin Length #2, Temporary Ventricular Sens Setting mV, Temporary Venticular Stim Threshold mA, Temperature Celsius, Temporary Venticular Sens Threshold mV, Troponin-T, Total Bilirubin, Mixed Venous O2% Sat, PeCO2
Pneumonia 9
ELWI (PiCCO), Recruitment Duration, T Low (APRV), CO (PiCCO), HDL, Cholesterol, SV (Arterial), Impaired Skin Width #5, LDL measured
Pneumothorax 10
HDL, Impaired Skin Width #3, Cardiac Output (thermodilution), Temporary Venticular Stim Threshold mA, TCO2 (calc) Venous, Total Bilirubin, Tidal Volume (set), Venous CO2 Pressure, Differential-Monos, Absolute Count - Eos

### v2

@@ Appendix 0.A Appendix
Atelectasis 10 CO (Arterial), HDL, Cholesterol, ELWI (PiCCO), T Low (APRV), GEDI (PiCCO), LDL measured, T High (APRV), LDL calculated, Serum Osmolality
Cardiomegaly 10 BiPap bpm (S/T -Back up), LDL measured, ELWI (PiCCO), D-Dimer, Impaired Skin Length #5, Impaired Skin Width #5, Uric Acid, GEDI (PiCCO), QTc, Cholesterol
Consolidation 9 Manual Blood Pressure Diastolic Right, Manual Blood Pressure Systolic Right, ELWI (PiCCO), GEDI (PiCCO), CFI (PiCCO), Manual Blood Pressure Diastolic Left, Negative Insp. Force, Cholesterol, PCA basal rate (mL/hour)
Edema 10 SV (Arterial), CO (Arterial), ELWI (PiCCO), CFI (PiCCO), GEDI (PiCCO), Bladder Scan Estimate, SVV (Arterial), Gentamicin (Random), LDL measured, BiPap bpm (S/T -Back up)
Enlarged Cardiomediastinum 5 ELWI (PiCCO), GEDI (PiCCO), SVV (Arterial), D-Dimer, RCexp (Measured Time Constant)
Fracture 10 Absolute Count - Monos, CK-MB, Absolute Count - Neuts, Troponin-T, CO2 production, Differential-Bands, Vti High, Absolute Count - Lymphs, Chloride (whole blood), Total Bilirubin
Lung Lesion 10 Temporary Ventricular Sens Setting mV, Temporary Venticular Sens Threshold mV, PCV Level, Absolute Count - Neuts, GI #1 Tube Mark (CM), Temporary Pacemaker Rate, Glucose (whole blood), Ionized Calcium, Total Bilirubin, Absolute Count - Eos
Lung Opacity 6 Cardiac Output (thermodilution), Bladder Scan Estimate, Ammonia, Serum Osmolality, PBP (Prefilter) Replacement Rate, Current Goal
Pleural Effusion 10 SV (Arterial), CO (Arterial), ELWI (PiCCO), Permanent Pacemaker Rate, GEDI (PiCCO), Gentamicin (Random), SVV (Arterial), Arctic Sun/Alsius Temp #2 C, Feeding Weight, Arctic Sun/Alsius Temp #1 C
Pleural Other 10 PCV Level, Impaired Skin Length #2, Temporary Ventricular Sens Setting mV, Temporary Venticular Stim Threshold mA, Temperature Celsius, Temporary Venticular Sens Threshold mV, Troponin-T, Total Bilirubin, Mixed Venous O2% Sat, PeCO2
Pneumonia 9 ELWI (PiCCO), Recruitment Duration, T Low (APRV), CO (PiCCO), HDL, Cholesterol, SV (Arterial), Impaired Skin Width #5, LDL measured
Pneumothorax 10 HDL, Impaired Skin Width #3, Cardiac Output (thermodilution), Temporary Venticular Stim Threshold mA, TCO2 (calc) Venous, Total Bilirubin, Tidal Volume (set), Venous CO2 Pressure, Differential-Monos, Absolute Count - Eos